OpenStudy (anonymous):

The difference of two numbers is 12. Two fifths of the greater number is six more than one third of the lesser number. Find both numbers. The difference of two numbers is 12. Two fifths of the greater number is six more than one third of the lesser number. Find both numbers. @Mathematics

OpenStudy (hoblos):

x-y=12 2x/5 = y/3 + 6 by solving the system we get: x=30 y=18

OpenStudy (aroub):

The first number: x The second number:x+12 2/5(x+12)=6+x/3 2/5x+24/5=6 +x/3 15(2/5x+24/5)=15(6+x/3) 6x+72=90+5x 6x-5x=-72+90 x= 18 the two numbers are: 18, 30 :)

OpenStudy (anonymous):

thanks aroub, clearly explained, i wish i can give you more than 1 metal

OpenStudy (aroub):

No problem! great, lol thank you! :D

OpenStudy (anonymous):

let the first number = x since the difference of both the numbers is 12 => the second number = x+12 now the greater number is obviously 'x+12' => 2/5(greater number)=6 more than the 1/3(smaller number) => 2/5(x+12)=6+x/3 => x=18 hence the numbers are 18 and 30 (the other number = x+12)